A local department store sells carpets in 3 sizes. Each carpet comes in 2 different qualities. One of the sizes comes in 6 colors. The other sizes come in 2 colors. How many choices of carpet are there?
Answer by stanbon(75887)

Let the sizes be A, B, and C
A comes in 2 qualities
Each quality has 6 different colors.
1*2*6=12 choices for A carpet
B comes in 2 qualities
Each quality has 2 colors
1*2*2=4 choices for B carpet
C is the same as B
1*2*2=4 choices for C carpet
Total number of choices = 20
You could see all this by constructing a tree diagram.
Cheers,
Stan H.